# PreAfford: Universal Affordance-Based Pre-Grasping for Diverse Objects and Environments

### v2

@@ Abstract
Robotic manipulation with two-finger grippers is challenged by objects lacking distinct graspable features. Traditional pre-grasping methods, which typically involve repositioning objects or utilizing external aids like table edges, are limited in their adaptability across different object categories and environments. To overcome these limitations, we introduce PreAfford, a novel pre-grasping planning framework that incorporates a point-level affordance representation and a relay training approach. Our method significantly improves adaptability, allowing effective manipulation across a wide range of environments and object types. When evaluated on the ShapeNet-v2 dataset, PreAfford not only enhances grasping success rates by 69% but also demonstrates its practicality through successful real-world experiments. These improvements highlight PreAfford’s potential to redefine standards for robotic handling of complex manipulation tasks in diverse settings.
@@ I Introduction
TABLE I: Comparisons with prior pre-grasping methods. This table offers a concise comparison of various pre-grasping studies, highlighting their key features and limitations. The last three columns evaluate each study’s adaptability across different object categories and environments, compatibility with existing grasping pipelines via a defined procedure to bypass the pre-grasping step, and deployability in real-world experiments without the need for custom shape representations [1] or identical environments to those used in simulators [2]. Columns labeled A, C, and D represent adaptability, compatibility, and deployability, respectively.
@@
We introduce the PreAfford, which effectively addresses the above limitations. The adaptability of PreAfford arises from a novel relay training paradigm within a dual-module framework. This setup generates a robust reward function based on the improvement in the likelihood of grasping success, which is consistently evaluated by the grasping module. For deployability and real-world adaptability, PreAfford employs a point-level affordance visual representation [15, 14, 16] that relies solely on RGB-D data. Additionally, the PreAfford integrates a pre-grasping necessity check at the start of the inference process, enabling direct grasping of objects that are straightforward to handle, thereby enhancing compatibility and operational efficiency.
@@
  • A novel, adaptive, and deployable pre-grasping framework that is also compatible with easy-to-grasp objects.
@@ II-A Pre-Grasping Tasks
The critical element is the formulation of the reward function for pre-grasping manipulation. Defining the graspability of an object within a specific environmental setup is challenging, leading to two main approaches for constructing the reward function. The first method involves manually programming the reward, using strategies like pre-defined goal regions [1] or specific transitions in the pose of the object and gripper [4, 13]. While effective within their specific experimental setups, these methods often lack flexibility to adapt to new object categories and environments. The second approach uses the output from pre-trained neural networks to estimate the graspability at a given pose [6, 2, 3]. This method typically relies on simulator-derived prior knowledge, such as signed distance functions [6] or precise object and environment geometries [2], which poses challenges for real-world application due to the necessity for a deep understanding of the scene [23, 24, 25, 26].
Furthermore, our findings indicate that many systems neglect the option to bypass the pre-grasping procedure for objects that can be directly grasped with a high confidence of success [6, 2, 3]. This oversight renders the pre-grasping framework redundant for the majority of daily graspable objects and unnecessarily increases the complexity and cost of robotic control. To date, no previous research has successfully integrated adaptability across various object-environment configurations, compatibility with directly graspable situations, and deployability in real-world scenarios in a single framework.
@@ II-B Point-Level Affordance for Robotic Manipulation
In our research, we expand the application of point-level affordance to guide pre-grasping manipulations, taking into account various geometries and environmental contexts. This advancement shows that point-level affordance can manage more complex scenarios with high accuracy and strong generalization potential.
@@ III-A Preliminaries
The primary goal of pre-grasping manipulation is to modify the poses of objects to increase the likelihood of a successful grasp. Our approach particularly focuses on the exploitation of environmental features.
@@
where $\vec{p}_{1}=(x_{1},y_{1},z_{1})$ and $\Delta\vec{x}_{1}=(\Delta x_{1},\Delta y_{1})$ denote the contact point location and horizontal displacement, respectively. The direction of pushing is by default set to be horizontal, and the gripper remains closed during this operation.
@@
where $p_{2}=(x_{2},y_{2},z_{2})$ denotes the contact point and $\vec{\theta}_{2}=(\alpha,\beta,\gamma)$ the Euler angles of the grasp orientation. When the gripper closes, it moves vertically by $\Delta z_{2}$ and maintains its grip for a duration $\Delta_{t}$. A grasp is labeled as successful if the object is lifted with a vertical displacement exceeding $\theta_{v}$ and the rotation $\sqrt{\alpha^{2}+\beta^{2}+\gamma^{2}}$ stays within $\theta_{r2}$.
@@ III-B Overview of the PreAfford
Figure 2: The framework of PreAfford. The framework consists of two main modules, each incorporating three networks: an affordance network, a proposal network, and a critic network. These networks respectively handle tasks of choosing the contact point, generating a proposal, and evaluating the proposal. PointNet++ (PN++) and MLP are employed to process point clouds and facilitate decision-making. During the inference phase, both modules collaborate to develop strategies for pre-grasping and grasping. In contrast, during the training phase, the grasping module generates rewards that are used to train the pre-grasping module, a process we refer to as relay.
@@
The training and inference processes operate in opposite directions, a concept we refer to as relay. The offline training dataset is collected from simulations. A grasping network, which judges the likelihood of a grasp’s success, is trained first to generate labels for training the pre-grasping module. Subsequently, the pre-grasping network is trained using both the simulation data and the outputs from the grasping network. During the prediction phase, the two modules form a closed-loop control system. The pre-grasping module adjusts the object’s pose until it is deemed suitable for the grasping module to apply a grasp.
@@ III-C Module Structure
A network $\mathcal{C}_{1}$ evaluates the efficacy of a pre-grasping operation by assigning a score $\mathcal{C}_{1}((p_{1},\Delta\vec{x}_{1})\mid\mathcal{O}_{1})\in\mathbb{R}$. This score considers the pair $(p_{1},\Delta\vec{x}_{1})$ proposed by $\mathcal{A}_{1}$ and $\mathcal{P}_{1}$. Similarly, $\mathcal{C}_{2}$ assesses the likelihood of success for a grasping operation, providing a score $\mathcal{C}_{2}(p_{2},\vec{\theta}_{2}\mid\mathcal{O}_{2})\in[0,1]$. Both critic networks utilize a single-layer MLP that processes the input features $f_{o}$, $f_{p_{i}}$, and $f_{M_{i}}$.
@@ III-D Inference
This procedure modifies the object’s pose within the environment, resulting in a new point cloud $\mathcal{O}_{2}$, which then serves as the input for the subsequent grasping module.
@@
Following the pre-grasping step, as discussed in Sec. III-D, the object is then ready to be grasped. The final selection of the manipulation technique is determined by the highest critic score. This score is produced by evaluating the combined output of the affordance network $\mathcal{A}_{2}$ and the proposal network $\mathcal{P}_{2}$. The manipulation with the highest score is then executed to grasp the object. This ensures that the chosen approach is optimal based on the current state of the object within the environment.
@@ III-E Training and Losses
Figure 4: Qualitative Results: Demonstrates training and testing categories in four scenarios—edge, slot, slope, and wall. Affordance maps highlight effective interaction areas, showing PreAfford ’s capability to devise suitable pre-grasping and grasping strategies for various object categories and scenes, including both seen and unseen objects.
@@ III-F Data Collection
Both datasets positioned the camera 3–5 meters away, oriented towards the object to simulate realistic robotic vision constraints. Each dataset included 10,000 successful and 30,000 unsuccessful interactions, emphasizing the need to account for negative outcomes in robotic tasks.
@@ IV-A Simulation Settings and Datasets
We tested PreAfford on 5 seen and 4 unseen categories of hard-to-grasp objects across 5 scenes to demonstrate its adaptability. These scenes include four with a single environmental feature (edge, wall, slope, and slot) and one novel scene (shown in Fig. 3) that contains all four features simultaneously, which is constructed to test adaptability to complex and unseen environments (labeled as Multi). We conducted 1,000 tests on each object-environment pair and calculated the mean success ratio. Other experimental setups are consistent with those described in Sec. III-F.
@@ IV-B Simulation Evaluation Metrics and Baselines
Figure 5: Real world pre-grasping manipulations with affordance maps. Red areas in the maps indicate optimal pushing locations. Point clouds are captured by Femto Bolt.
@@ IV-C Simulation Results and Analysis
TABLE II: Comparison with baselines. Pre-grasping increases grasping success rates by 52.9%. A closed-loop strategy further enhances this improvement by 16.4% across all categories.
@@ IV-D Real-world Experiment Setup
We implemented our algorithm on the AIRBOT Play robotic arm, a compact six-degree-of-freedom manipulator. For capturing RGB-D data, we used an ORBBEC Femto Bolt camera. The end effector of the system is an INSPIRE-ROBOTS EG2-4C gripper.
@@ V Conclusion and Limitation
In this paper, we introduced a novel two-stage affordance learning framework that excels in adaptability across various object-environment configurations, shows compatibility with graspable objects, and proves deployable through rigorous real-world testing. This framework has been validated in both simulated and real-world settings, underscoring its effectiveness. However, we acknowledge that our method still faces challenges, particularly in handling objects with highly irregular shapes or extremely dynamic environments. Future work will focus on enhancing the robustness and flexibility of our approach to address these issues.

### v3

@@ Abstract
Robotic manipulation with two-finger grippers is challenged by objects lacking distinct graspable features. Traditional pre-grasping methods, which typically involve repositioning objects or utilizing external aids like table edges, are limited in their adaptability across different object categories and environments. To overcome these limitations, we introduce PreAfford, a novel pre-grasping planning framework incorporating a point-level affordance representation and a relay training approach. Our method significantly improves adaptability, allowing effective manipulation across a wide range of environments and object types. When evaluated on the ShapeNet-v2 dataset, PreAfford not only enhances grasping success rates by 69% but also demonstrates its practicality through successful real-world experiments. These improvements highlight PreAfford’s potential to redefine standards for robotic handling of complex manipulation tasks in diverse settings.
@@ I Introduction
TABLE I: Comparisons with prior pre-grasping methods. This table offers a concise comparison of various pre-grasping studies, highlighting their key features and limitations. The last three columns evaluate each study’s adaptability across different object categories and environments, compatibility with existing grasping pipelines via a defined procedure to bypass the pre-grasping step, and deployability in real-world experiments without the need for custom shape representations [1] or identical environments to those used in simulators [2]. Columns A, C, and D represent adaptability, compatibility, and deployability, respectively.
@@
We introduce the PreAfford, which effectively addresses the above limitations. The adaptability of PreAfford arises from a novel relay training paradigm within a dual-module framework. This setup generates a robust reward function to improve the likelihood of grasping success, which is consistently evaluated by the grasping module. For deployability and real-world adaptability, PreAfford employs a point-level affordance visual representation [15, 14, 16] that relies solely on RGB-D data. Additionally, the PreAfford integrates a pre-grasping necessity check at the start of the inference process, enabling direct grasping of objects that are straightforward to handle, thereby enhancing compatibility and operational efficiency.
@@
  • A novel, adaptive, and deployable pre-grasping framework, compatible with easy-to-grasp objects.
@@ II-A Pre-Grasping Tasks
The critical element is the formulation of the reward function for pre-grasping manipulation. Defining the graspability of an object within a specific environmental setup is challenging, leading to two main approaches for constructing the reward function. The first method involves manually programming the reward, using strategies like pre-defined goal regions [1] or specific transitions in the pose of the object and gripper [4, 13]. Although effective within their specific experimental setups, these methods often lack the flexibility to adapt to new object categories and environments. The second approach uses the output from pre-trained neural networks to estimate the graspability at a given pose [6, 2, 3]. This method typically relies on simulator-derived prior knowledge, such as signed distance functions [6] or precise object and environment geometries [2], which poses challenges for real-world application due to the necessity for a deep understanding of the scene [23, 24, 25, 26].
Furthermore, our findings indicate that many systems neglect the option to bypass the pre-grasping procedure for objects that can be directly grasped with a high confidence of success [6, 2, 3]. This oversight renders the pre-grasping framework redundant for most daily graspable objects and unnecessarily increases the complexity and cost of robotic control. To date, no previous research has successfully integrated adaptability across various object-environment configurations, compatibility with directly graspable situations, and deployability in real-world scenarios in a single framework.
@@ II-B Point-Level Affordance for Robotic Manipulation
In our research, we expand the application of point-level affordance to guide pre-grasping manipulations, considering various geometries and environmental contexts. This advancement shows that point-level affordance can manage more complex scenarios with high accuracy and strong generalization potential.
@@ III-A Preliminaries
The primary goal of pre-grasping manipulation is to modify object poses to increase the likelihood of a successful grasp. Our approach particularly focuses on the exploitation of environmental features.
@@
where $\vec{p}_{1}=(x_{1},y_{1},z_{1})$ and $\Delta\vec{x}_{1}=(\Delta x_{1},\Delta y_{1})$ denote the contact point location and horizontal displacement, respectively. By default, the direction of pushing is set to horizontal, and the gripper remains closed during this operation.
@@
where $p_{2}=(x_{2},y_{2},z_{2})$ denotes the contact point and $\vec{\theta}_{2}=(\alpha,\beta,\gamma)$ the Euler angles of the grasp orientation. When the gripper closes, it moves vertically by $\Delta z_{2}$ and maintains its grip for a duration $\Delta_{t}$. A grasp is labeled as successful if the object is lifted with a vertical displacement exceeding $\theta_{v}$ and the rotation $\sqrt{\alpha^{2}+\beta^{2}+\gamma^{2}}$ remains within $\theta_{r2}$.
@@ III-B Overview of the PreAfford
Figure 2: The framework of PreAfford. The framework consists of two main modules, each incorporating three networks: an affordance network, a proposal network, and a critic network. These networks respectively handle tasks of choosing the contact point, generating a proposal, and evaluating the proposal. PointNet++ (PN++) and MLP are employed to process point clouds and facilitate decision-making. During the inference phase, both modules collaborate to develop strategies for pre-grasping and grasping. In contrast, during the training phase, the grasping module generates rewards for training the pre-grasping module, a process we refer to as relay.
@@
Training and inference processes operate in opposite directions, a concept we refer to as relay. The offline training dataset is collected from simulations. A grasping network, which judges the likelihood of a grasp’s success, is trained first to generate labels for training the pre-grasping module. Subsequently, the pre-grasping network is trained using both the simulation data and the outputs from the grasping network. During the prediction phase, the two modules form a closed-loop control system. The pre-grasping module adjusts the object’s pose until it is deemed suitable for the grasping module to apply a grasp.
@@ III-C Module Structure
A network $\mathcal{C}_{1}$ evaluates the efficacy of a pre-grasping operation by assigning a score $\mathcal{C}_{1}((p_{1},\Delta\vec{x}_{1})\mid\mathcal{O}_{1})\in\mathbb{R}$. This score considers the pair $(p_{1},\Delta\vec{x}_{1})$ proposed by $\mathcal{A}_{1}$ and $\mathcal{P}_{1}$. Similarly, $\mathcal{C}_{2}$ assesses the likelihood of success for a grasping operation, providing a score $\mathcal{C}_{2}(p_{2},\vec{\theta}_{2}\mid\mathcal{O}_{2})\in[0,1]$. Both critic networks utilize an MLP to process the input features $f_{o}$, $f_{p_{i}}$, and $f_{M_{i}}$.
@@ III-D Inference
This procedure modifies the object’s pose within the environment, resulting in a new point cloud $\mathcal{O}_{2}$, which serves as the input for the subsequent grasping module.
@@
Following the pre-grasping step, as discussed in Sec. III-D, the object is ready to be grasped. The final selection of the manipulation technique is determined by the highest critic score. This score is produced by evaluating the combined output of the affordance network $\mathcal{A}_{2}$ and the proposal network $\mathcal{P}_{2}$. The manipulation with the highest score is then executed to grasp the object. This ensures that the chosen approach is optimal based on the current state of the object within the environment.
@@ III-E Training and Losses
Figure 4: Qualitative Results. We demonstrate pre-grasping manipulation on training and testing categories in four scenarios—edge, slot, slope, and wall. Affordance maps highlight effective interaction areas, showing PreAfford’s capability to devise suitable pre-grasping and grasping strategies for various object categories and scenes, including both seen and unseen objects.
@@ III-F Data Collection
Both datasets positioned the camera 3–5 meters away, oriented towards the object to simulate realistic robotic vision constraints. Each dataset included 10,000 successful and 30,000 unsuccessful interactions. The data-collection process takes about 8 days on 256 AMD EPYC 7742 processors.
@@ IV-A Simulation Settings and Datasets
We tested PreAfford on 5 seen (hat, laptop, scissors, keyboard, and tablet) and 4 unseen (phone, pliers, switch, and window) categories of hard-to-grasp objects across 5 scenes. These scenes include four with a single environmental feature (edge, wall, slope, and slot) and one novel scene (shown in Fig. 3) that contains all four features simultaneously, which is constructed to test adaptability to complex and unseen environments. We conducted 1,000 tests on each object-environment pair and calculated the mean success ratio.
@@ IV-B Simulation Evaluation Metrics and Baselines
Figure 5: Real-world pre-grasping manipulations with affordance maps. Red areas in the maps indicate optimal pushing locations. Point clouds are captured by Femto Bolt.
@@ IV-C Simulation Results and Analysis
TABLE II: Success rates of grasping manipulation in percentage. Pre-grasping increases grasping success rates by 52.9%. A closed-loop strategy further enhances this improvement by 16.4% across all categories.
@@ IV-D Real-world Experiment Setup
We implemented our algorithm on the AIRBOT Play robotic arm, a compact six-degree-of-freedom manipulator. To capture RGB-D data, we used an ORBBEC Femto Bolt camera. The system’s end effector is an INSPIRE-ROBOTS EG2-4C gripper.
@@ V Conclusion and Limitation
In this paper, we introduced a novel two-stage affordance learning framework that excels in adaptability across various object-environment configurations, shows compatibility with graspable objects, and proves deployable through rigorous real-world testing. This framework has been validated in simulation and real-world settings, underscoring its effectiveness. However, our method still faces challenges, particularly in handling objects with highly irregular shapes or extremely dynamic environments. Future work will focus on enhancing the robustness and flexibility of our approach to address these issues.